OpenStudy (anonymous):

Use substitution to solve. 2x+3y=1 -x+(1/3)y=5

OpenStudy (mathstudent55):

1. Multiply entire 2n eq by 3 to get rid of fraction: -3x + y = 15 Now solve it for y: y = 15 + 3x 2. Now substitute what y is equal to (15 + 3x), where you see y in the first original equation: 2x + 3(15 + 3x) = 1 and solve for x. Then once you have x, substitute that into x in the first or second original equation.